# A CAV-based perimeter-free regional traffic control strategy utilizing existing parking infrastructure

### v2

@@ Abstract
This paper proposes a novel perimeter-free regional traffic management strategy for traffic networks under a connected and autonomous vehicle (CAV) environment. The proposed strategy requires CAVs, especially those with long remaining travel distances, to temporarily wait at nearby parking facilities when the network is congested. After a designated holding time, these CAVs are allowed to re-enter the network. Doing so helps reduce congestion and improve overall operational efficiency. Unlike traditional perimeter control approaches that restrict inflows to congested regions, the proposed holding strategy leverages existing parking infrastructure to temporarily hold vehicles in a way that partially avoids local queue accumulation issues. The proposed method can be easily integrated with existing signal control methods and retains the maximum stability property of the original traffic signal control methods. Simulation results show that the proposed strategy not only reduces travel time for vehicles that are not held, but can also reduce travel times for some of the held vehicles as well, which serves as another key merit of the proposed approach. Compared to the two benchmark perimeter control algorithms, the proposed strategy is more robust against demand patterns and generates stronger improvements in the operational efficiency. Importantly, since the proposed strategy requires existing parking infrastructure, its performance has been demonstrated under various configurations of parking locations and capacities. Particularly, it is demonstrated that the utilization of the parking facility consistently improves overall traffic efficiency, regardless of the facility’s size. Lastly, the proposed strategy is shown to be beneficial in a partial CAV environment where only a subset of vehicles are available for holding. The results offer insights into the potential applicability of the proposed strategy to human-driven vehicles, provided that drivers are assured their temporary wait could possibly lead to a reduction in their own travel time.
@@ 1 Introduction
Despite substantial research effort in the area of traffic flow management, traffic congestion remains a demanding issue due to the continuous increase in vehicular traffic. In urban networks, traffic signals are one of the most cost effective means to manage traffic flows at intersections, which serve as the primary recurring bottlenecks in urban areas. Based on the spatial scale that is considered when making signal timing decisions, traffic signal control methods can be classified as isolated intersection control (e.g., Wunderlich et al. (2008); Yang et al. (2016); Yu et al. (2018); Wan and Hwang (2018)), coordinated signal control along arterial streets (e.g., Bazzan (2005); Zhang et al. (2015); Li and Ban (2020)), and network-wide traffic signal control (e.g., Mirchandani and Head (2001); Arel et al. (2010); Varaiya (2013); Chen et al. (2020)).
@@
Compared to isolated and coordinated arterial level control, a well designed network-wide traffic signal control algorithm can be more beneficial for the operational efficiency of the entire network. However, due to the complexity and computational burden resulting from interdependence between the large number of elements, network-wide control remains a challenging task. Two main categories of network-wide traffic signal control exist based on how decisions at individual intersections are made relative to one another: centralized control (e.g., Lowrie (1990); Robertson and Bretherton (1991); Park et al. (1999); Li et al. (2016); Liang et al. (2019)) and decentralized control (e.g., Lämmer and Helbing (2008); Varaiya (2013); Kouvelas et al. (2011); Rinaldi et al. (2016); Sha and Chow (2019); Chen et al. (2020)). Centralized control methods usually rely on large-scale optimization programs in which signal timings at multiple intersections are determined jointly considering their influence on each other. Although this method can potentially achieve global optimal solutions, they are usually not scalable due to the computational complexity involved. On the contrary, signal timings at each intersection derived from decentralized control methods are independent of other intersections; however, the computational speed of this strategy is usually higher than the centralized control methods. Thanks to the computational efficiency, which is a crucial factor for real-time traffic signal control, more research effort has been shifted to this method in the past decade.
@@
In general, both centralized and decentralized network-wide traffic signal control strategies require estimates of traffic state, such as queue length, at local intersections, which is difficulty to satisfy for most intersections. In addition, most of those algorithms do not consider the relationship between network-level traffic conditions and local signal timings. As a result, they cannot prevent or tackle regional congestion effectively. Perimeter control, which is a method derived based on the theory of Macroscopic Fundamental Diagram (MFD) (Godfrey, 1969; Daganzo, 2007; Geroliminis and Daganzo, 2008), is an effective strategy to tackle both issues. A typical MFD reflects the uni-modal and concave relationship between the average operational efficiency, represented by average flow or vehicle exit rate, and the average use, represented by average density or vehicle accumulation, in a specific region of a network. In general, there exists a critical value for the average use of the network at which the maximum operational efficiency at a network-level can be achieved. Under high demand conditions, the network enters the congested domain once the average use exceeds this critical value, leading to a reduction in the operational efficiency. When traffic demand for the target region is high enough to lead congestion, perimeter control limits the inflows to the target region by adding restrictions at the perimeter of the region to maintain the accumulation or density inside the target region around the critical value, so that the operational efficiency can be near the maximum. Compared to the network-wide traffic signal control algorithms mentioned in the previous paragraph, perimeter metering control does not require the traffic state knowledge at all intersections, and it can result in improvements to overall network-level operations.
@@
Various perimeter control algorithms have been developed in the literature. For example, Bang-Bang control (Daganzo, 2007) allows no vehicle to enter the region whenever the accumulation exceeds the critical value and removes the restriction only once the density drops below the critical value. This strategy has been demonstrated to be an optimal control strategy for abstract systems in which the effect of queue accumulation is not considered. Another widely used perimeter control strategy is proportion-integral-type (PI) feedback control (Keyvan-Ekbatani et al., 2012; Ramezani et al., 2015; Haddad and Shraiber, 2014). Unlike Bang-Bang control, the restriction is adjusted based on the inflows in the previous time step, the change of the average density in the previous time step, and the distance between the current density and the critical density. Therefore, the restriction from this strategy is intensified or loosened more smoothly than Bang-Bang control. Reinforcement learning-based approaches (Chu et al., 2016; Ni and Cassidy, 2019; Zhou and Gayah, 2021; Chen et al., 2022; Zhou and Gayah, 2023) have also recently been developed and shown to provide reasonable performances compared to MFD theory-based models.
@@
Algorithms that combine perimeter control and decentralized control algorithms have been proposed to further enhance the efficiency of urban transportation systems. For example, Tsitsokas et al. (2023) developed a bi-level control framework to combine PI feedback control and Max Pressure (MP) algorithm (Tassiulas and Ephremides, 1990; Varaiya, 2013), which is a popular decentralized traffic signal control algorithm, and unveiled that the proposed approach can further enhance the efficiency compared to a purely MP controlled network. Su et al. (2023) developed a similar approach to integrate reinforcement learning perimeter control algorithm and MP control. In both studies, perimeter control and MP are separately implemented at perimeter intersections and other intersections, respectively. More recently, Liu and Gayah (2024) proposed an algorithm, called N-MP, to incorporate perimeter control mechanism directly within the MP control structure. The N-MP algorithm still follows the decentralized control architecture of traditional MP algorithms and can tackle the heterogeneous vehicle distribution inside the protected region.
@@
Despite the promising performance of perimeter control algorithms, there exist common drawbacks that can limit their potential. For one, most perimeter control approaches require the identification of perimeter intersections, which are both pre-determined and fixed. However, as pointed out in (Li et al., 2021; Doig et al., 2024), a stationary perimeter can impede the full potential of perimeter control when traffic conditions vary in both the temporal and spatial dimensions, which is not an unusual phenomenon for congested regions. More importantly, all perimeter control algorithms lead to temporary queue accumulation at the external boundary of the perimeter. These queues can not only reduce the operation efficiency by blocking destinations at the external region, but they can also keep vehicles from exiting the protected region if spillover occurs and might also contribute to gridlock-like conditions near the perimeter.
@@
Triggered by these drawbacks, this paper considers a perimeter-free control strategy in a connected and autonomous vehicle (CAV) environment to deal with regional congestion. The proposed method assumes that CAVs are able to receive and will comply with requests to temporarily exit the network and wait at nearby parking facilities in order to improve overall network efficiency. In particular, when the network is congested, a subset of CAVs - specifically those with significant remaining travel distances - are requested to use nearby parking spaces to wait. Doing so can help mitigate congestion and reduce travel time for the vehicles with shorter remaining travel distances, allowing them to exit the network more quickly and reduce overall congestion levels. After a certain time, the held CAVs are allowed to re-enter the network to complete their trips. The proposed strategy harnesses both the control flexibility from CAVs and the available parking infrastructure to reduce the negative impact of vehicle accumulation at perimeter boundaries and further improve overall operational efficiency. Note, while studies have explored various aspects of CAV parking, including motion planning (Hsieh and Ozguner, 2008; Wang and Zhu, 2013), parking lots layout design (Nourinejad et al., 2018; Yan et al., 2022), parking pricing (Wang et al., 2021; Fulman and Benenson, 2018), to the best of our knowledge, no research has examined the potential benefits of utilizing existing parking facilities for CAVs to improve operational efficiency.
@@
Any traffic signal control algorithm can be used as the base control for the proposed strategy at intersections within the entire network. In this paper, we evaluate the performance of the proposed strategy integrated with the queue-based MP control developed by Varaiya (2013), using microscopic traffic simulation. To demonstrate its effectiveness as a network-level traffic management strategy, we compare it against existing perimeter control strategies, which are among the most effective methods for mitigating regional congestion. Simulation results show that the proposed approach can outperform the classical Bang-Bang control strategy and N-MP under various traffic conditions, even under partial CAV environments in which only a subset of vehicles are eligible to be held. Interestingly, the proposed algorithm can reduce travel time for both held and non-held vehicles. Moreover, we proved that if the base signal control has the maximum stability property, it still holds after implementing the proposed holding strategy.
@@ 2 Methodology
We assume that all intersections in the network are controlled under a base signal control algorithm. We consider both fully CAV environments, where all vehicles comply with the control policy and temporarily park when requested, and mixed humand driven vehicle (HDV)-CAV environments, where only CAVs adhere to the parking requests. According to MFD theory, there exists a critical value for the average density, $\rho_{cr}$, at which the network efficiency is maximized. In the proposed algorithm, once the average density, $\rho(t)$, exceeds $\rho_{cr}$, the CAVs with a remaining travel distance longer than a threshold value, $\phi$, are requested to park at a nearby parking space, if a space is available. After a holding time of $\tau$, the held CAVs will attempt to re-enter the network to finish their trips as soon as there exists available space on the road. Note only vehicles that are already passing the parking locations at the moment when $\rho(t)>\rho_{cr}$ will be held at the associated parking locations, and held vehicles will re-enter the network through the road from which they exit the network.
@@ 2.1 Traffic Dynamics
where $C_{i,j,k}(t)$ is the saturation flow of movement $(i,j,k)$ at time $t$. The exact number of departing CAVs and HDVs, denoted by $y_{i,j,k}^{\text{CAV,P}}(t)$ and $y_{i,j,k}^{\text{HDV}}(t)$, are dependent on the vehicle distribution on link $(i,j)$ and the vehicle transition models. For instance, under a First-In-First-Out (FIFO) principle, the present CAVs and HDVs occupying the first $y_{i,j,k}(t)$ positions in the queue will depart upon receiving a green time.
@@
$n^{\text{enter}}_{i,j,k}(t)=\sum_{v=1}^{x_{i,j,k}^{\text{CAV,H}}(t)}\mathbbm{1}(e_{v}(t)==\tau)$ (5)
Note that $e_{v}(t)$ accounts for the delay incurred due to both parking and re-entering the network. In other words, the waiting time includes the duration from when a CAV starts to enter an available parking spot to when it re-enters the roadway in the network. Further, while ignored in this section, the simulation results explicitly consider traffic impacts of vehicles exiting/reentering the network before/after their parking maneuvers.
@@
$n^{\text{hold}}_{i,j,k}(t)=\min\left(\mathbbm{1}_{[\rho_{\text{cr}},+\infty]}\rho(t)\sum_{v=1}^{x_{i,j,k}^{\text{CAV,P}}(t)+q^{\text{CAV,P}}_{i,j,k}(t)}\mathbbm{1}_{[\phi,+\infty]}(g_{v}(t)),K^{p}_{i,j,k}-x_{i,j,k}^{\text{CAV,H}}(t)\right)$ (6)
where $g_{v}(t)$ is the remaining travel distance of vehicle $v$ at time $t$, and $K^{p}_{i,j,k}$ is the total parking capacity associated with movement $(i,j,k)$, which serves as the upper bound of the number of held CAVs from movement $(i,j,k)$. Note this paper assumes that the layout of parking resources, including the locations and parking capacities, $K^{p}$’s, is known. The first term in Eq. (6) is the holding demand, and the second term is the available parking supply for movement $(i,j,k)$, respectively.

### v3

# A CAV-based perimeter-free regional traffic control utilizing existing parking infrastructure

@@ Abstract
This paper proposes a novel perimeter-free regional traffic management strategy for networks under a connected and autonomous vehicle (CAV) environment. The proposed strategy requires a subset of CAVs to temporarily wait at nearby parking facilities when the network is congested. After a designated holding time, these CAVs are allowed to re-enter the network. Doing so helps reduce congestion and improve overall operational efficiency. Unlike traditional perimeter control approaches, the proposed strategy leverages existing parking infrastructure to temporarily hold vehicles in a way that partially avoids local queue accumulation issues. Further, holding the vehicles with the longest remaining travel distances creates a self-reinforcing mechanism which helps reduce congestion more quickly than perimeter metering control. Simulation results show that the proposed strategy not only reduces travel time for vehicles that are not held, but can also reduce travel times for some of the held vehicles as well. Importantly, its performance has been demonstrated under various configurations of parking locations and capacities and CAV penetration rates.
@@ 1 Introduction
Despite substantial research effort in the area of traffic flow management, traffic congestion remains a demanding issue due to the continuous increase in vehicular traffic. In urban networks, traffic signals are one of the most cost effective means to manage traffic flows at intersections, which serve as the primary recurring bottlenecks in urban areas. Based on the spatial scale that is considered when making signal timing decisions, traffic signal control methods can be classified as isolated intersection control (e.g., [53, 57, 58, 47]), coordinated signal control along arterial streets (e.g., [3, 59, 22]), and network-wide traffic signal control (e.g., [31, 2, 46, 5]).
@@
Compared to isolated and coordinated arterial level control, a well designed network-wide traffic signal control algorithm can be more beneficial for the operational efficiency of the entire network. However, due to the complexity and computational burden resulting from interdependence between the large number of elements, network-wide control remains a challenging task. Two main categories of network-wide traffic signal control exist based on how decisions at individual intersections are made relative to one another: centralized control (e.g., [29, 38, 35, 21, 24]) and decentralized control (e.g., [18, 46, 16, 37, 39, 5]). Centralized control methods usually rely on large-scale optimization programs in which signal timings at multiple intersections are determined jointly considering their influence on each other. Although this method can potentially achieve global optimal solutions, they are usually not scalable due to the computational complexity involved. On the contrary, signal timings at each intersection derived from decentralized control methods are independent of other intersections; however, the computational speed of this strategy is usually higher than the centralized control methods. Thanks to the computational efficiency, which is a crucial factor for real-time traffic signal control, more research effort has been shifted to this method in the past decade.
@@
In general, both centralized and decentralized network-wide traffic signal control strategies require estimates of traffic state, such as queue length, at local intersections, which is difficulty to satisfy for most intersections. In addition, most of those algorithms do not consider the relationship between network-level traffic conditions and local signal timings. As a result, they cannot prevent or tackle regional congestion effectively. Perimeter control, which is a method derived based on the theory of Macroscopic Fundamental Diagram (MFD) [12, 7, 11], is an effective strategy to tackle both issues. A typical MFD reflects the uni-modal and concave relationship between the average operational efficiency, represented by average flow or vehicle exit rate, and the average use, represented by average density or vehicle accumulation, in a specific region of a network. In general, there exists a critical value for the average use of the network at which the maximum operational efficiency at a network-level can be achieved. Under high demand conditions, the network enters the congested domain once the average use exceeds this critical value, leading to a reduction in the operational efficiency. When traffic demand for the target region is high enough to lead congestion, perimeter control limits the inflows to the target region by adding restrictions at the perimeter of the region to maintain the accumulation or density inside the target region around the critical value, so that the operational efficiency can be near the maximum. Compared to the network-wide traffic signal control algorithms mentioned in the previous paragraph, perimeter metering control does not require the traffic state knowledge at all intersections, and it can result in improvements to overall network-level operations.
@@
Various perimeter control algorithms have been developed in the literature. For example, Bang-Bang control [7] allows no vehicle to enter the region whenever the accumulation exceeds the critical value and removes the restriction only once the density drops below the critical value. This strategy has been demonstrated to be an optimal control strategy for abstract systems in which the effect of queue accumulation is not considered. Another widely used perimeter control strategy is proportion-integral-type (PI) feedback control [15, 36, 13]. Unlike Bang-Bang control, the restriction is adjusted based on the inflows in the previous time step, the change of the average density in the previous time step, and the distance between the current density and the critical density. Therefore, the restriction from this strategy is intensified or loosened more smoothly than Bang-Bang control. Reinforcement learning-based approaches [6, 32, 60, 4, 61] have also recently been developed and shown to provide reasonable performances compared to MFD theory-based models.
@@
Algorithms that combine perimeter control and decentralized control algorithms have been proposed to further enhance the efficiency of urban transportation systems. For example, [45] developed a bi-level control framework to combine PI feedback control and Max Pressure (MP) algorithm [44, 46], which is a popular decentralized traffic signal control algorithm, and unveiled that the proposed approach can further enhance the efficiency compared to a purely MP controlled network. [42] developed a similar approach to integrate reinforcement learning perimeter control algorithm and MP control. In both studies, perimeter control and MP are separately implemented at perimeter intersections and other intersections, respectively. More recently, [28] proposed an algorithm, called N-MP, to incorporate perimeter control mechanism directly within the MP control structure. The N-MP algorithm still follows the decentralized control architecture of traditional MP algorithms and can tackle the heterogeneous vehicle distribution inside the protected region.
@@
Despite the promising performance of perimeter control algorithms, there exist common drawbacks that can limit their potential. For one, most perimeter control approaches require the identification of perimeter intersections, which are both pre-determined and fixed. However, as pointed out in [23, 9], a stationary perimeter can impede the full potential of perimeter control when traffic conditions vary in both the temporal and spatial dimensions, which is not an unusual phenomenon for congested regions. More importantly, all perimeter control algorithms lead to temporary queue accumulation at the external boundary of the perimeter. These queues can not only reduce the operation efficiency by blocking destinations at the external region, but they can also keep vehicles from exiting the protected region if spillover occurs and might also contribute to high congestion and gridlock-like conditions near the perimeter.
@@
Triggered by these drawbacks, this paper considers a perimeter-free control strategy to deal with regional congestion. The proposed method assumes that a subset of vehicles are able to receive and will comply with requests to temporarily exit the network and wait at nearby parking facilities in order to improve overall network efficiency. In particular, when the network is congested, vehicles with significant remaining travel distances are requested to use nearby parking spaces to wait. Doing so can help mitigate congestion and reduce travel time, allowing them to exit the network more quickly and reduce overall congestion levels. After a certain time, the held vehicles are allowed to re-enter the network to complete their trips. Importantly, prioritizing long-distance trips for holding creates a self-reinforcing congestion mitigation mechanism. First, temporarily removing these vehicles directly decreases network accumulation. Second, because the remaining vehicles tend to have shorter remaining distances, they can complete their trips more quickly, which increases the vehicle exit rate and further accelerates congestion dissipation.
We assume for simplicity that the vehicles that are held are Connected and Autonomous Vehicles (CAVs). These CAVs are compliant with the holding requests for two reasons. First, many CAVs—such as delivery vehicles, empty cruising taxis, and fleet-operated vehicles—do not involve human passengers. These vehicles can fully comply with holding instructions without behavioral resistance. Second, even for occupied CAVs, the perception of holding from CAV passengers is different from that from drivers of HDVs. The former can continue working or engaging in other activities while the vehicle is being held, and they do not face the burden associated with re-entering traffic. As a result, their perceived inconvenience is expected to be lower, leading to higher expected compliance compared with human drivers. Note in this study, we assume that CAVs and human-driven vehicles (HDVs) share identical driving behavior parameters; the only distinction is that CAVs comply with holding instructions.
The proposed strategy harnesses both the control flexibility from CAVs and the available parking infrastructure to reduce the negative impact of vehicle accumulation at perimeter boundaries and further improve overall operational efficiency. The results offer insights for policymakers in creating and allocating parking facilities to support the implementation of this strategy. We would like to emphasize that the primary objective of this paper is to explore, for the first time, the benefits and feasibility of the proposed holding strategy as a novel traffic flow management approach, paving the way for relevant future research. Developing detailed theoretical models to enhance the strategy, including the optimization of critical parameters of the proposed method, is beyond the scope of this paper and will be addressed in future studies.
While previous studies have explored various aspects of CAV parking, including motion planning [14, 50], parking lot layout [34, 56], parking pricing [48, 10], the objective of this paper is to explore the benefits of a proposed holding strategy that utilizes parking facilities in a CAV environment to improve real-time traffic operational efficiency—an area that, to the best of our knowledge, has not yet been studied. Therefore, the literature review focuses on traffic flow management strategies, including traffic signal control and perimeter control, which are popular network-wide techniques for mitigating regional traffic congestion, rather than on parking-related research that addresses other perspectives. One relevant study [55] investigated parking provisions for e-hailing ride-sourcing vehicles (RVs), proposing a parking strategy for vacant RVs to reduce cruising vehicles in the network and thereby alleviate traffic congestion. Our objective, however, differs in that we focus on general private vehicles, including CAVs and potentially HDVs, which leads to different findings regarding the impact of parking provisions on traffic conditions. Overall, our results show that providing parking spaces during peak hours can help reduce congestion, whereas that study concluded that parking provisions for RVs are only necessary during non-peak hours, as cruising is minimal in peak periods.
@@
Any traffic signal control algorithm can be used as the base control for the proposed strategy at intersections within the entire network. In this paper, we evaluate the performance of the proposed strategy integrated with the queue-based MP control developed by [46], using microscopic traffic simulation. To demonstrate its effectiveness as a network-level traffic management strategy, we compare it against existing perimeter control strategies, which are among the most effective methods for mitigating regional congestion. Simulation results show that the proposed approach can outperform the classical Bang-Bang control strategy and N-MP under various traffic conditions, even under partial CAV environments in which only a subset of vehicles are eligible to be held. Interestingly, the proposed algorithm can reduce travel time for both held and non-held vehicles. Moreover, we proved that if the base signal control has the maximum stability property, it still holds after implementing the proposed holding strategy.
@@ 2 Methodology
We assume that all intersections in the network are controlled under a base signal control algorithm. [^1] Any traffic signal control algorithm can be used as the base control for the proposed strategy at intersections within the entire network. In this paper, we evaluate the performance of the proposed strategy integrated with the queue-based MP control developed by [46], using microscopic traffic simulation. We consider both fully CAV environments, where all vehicles comply with the control policy and temporarily park when requested, and mixed HDV-CAV environments, where only CAVs adhere to the parking requests. According to MFD theory, there exists a critical value for the average density, $\rho_{cr}$, at which the network efficiency is maximized. In the proposed algorithm, once the average density, $\rho(t)$, exceeds $\rho_{cr}$, the CAVs with a remaining travel distance longer than a threshold value, $\phi$, are requested to park at a nearby parking space, if a space is available. The decision to hold long-distance trips is justified both theoretically (see Section 2.2) and numerically (see Section 4.2.5). After a holding time of $\tau$, the held CAVs will attempt to re-enter the network to finish their trips as soon as there exists available space on the road. Note only vehicles that are already passing the parking locations at the moment when $\rho(t)>\rho_{cr}$ will be held at the associated parking locations, and held vehicles will re-enter the network through the road from which they exit the network. As a result, this strategy does not generate cruising or rerouting maneuvers for parking searching; therefore, the impacts associated with such maneuvers are not considered in this paper.
@@ 2.1 Traffic Dynamics
where $C_{i,j,k}(t)$ is the saturation flow of movement $(i,j,k)$ at time $t$. The exact number of departing CAVs and HDVs, denoted by $y_{i,j,k}^{\text{CAV,P}}(t)$ and $y_{i,j,k}^{\text{HDV}}(t)$, are dependent on the vehicle distribution on link $(i,j)$ and the vehicle transition models. For instance, under a First-In-First-Out (FIFO) principle, the present CAVs and HDVs occupying the first $y_{i,j,k}(t)$ positions in the queue will depart upon receiving a green time. However, our method does not rely on specific queue serving disciplines.
@@
$n^{\text{enter}}_{i,j,k}(t)=\sum_{v=1}^{x_{i,j,k}^{\text{CAV,H}}(t)}\mathbbm{1}(e_{v}(t)>=\tau)$ (5)
Note that $e_{v}(t)$ accounts for the delay incurred due to both parking and re-entering the network. In other words, the waiting time includes the duration from when a CAV exits the travel lanes to enter an available parking spot to when it re-enters the roadway in the network. Further, while ignored in this section, the simulation results explicitly consider traffic impacts of vehicles exiting/reentering the network before/after their parking maneuvers.
@@
$n^{\text{hold}}_{i,j,k}(t)=\min\left(\mathbbm{1}_{[\rho_{\text{cr}},+\infty]}\rho(t)\sum_{v=1}^{x_{i,j,k}^{\text{CAV,P}}(t)+q^{\text{CAV,P}}_{i,j,k}(t)}(1-\delta_{v})\mathbbm{1}_{[\phi,+\infty]}(g_{v}(t)),K^{p}_{i,j,k}-x_{i,j,k}^{\text{CAV,H}}(t)\right)$ (6)
where $\delta_{v}$ denotes a binary variable indicating whether vehicle $v$ has been held, taking the value 1 if held and 0 otherwise; $g_{v}(t)$ is the remaining travel distance of vehicle $v$ at time $t$, and $K^{p}_{i,j,k}$ is the total parking capacity associated with movement $(i,j,k)$, which serves as the upper bound of the number of held CAVs from movement $(i,j,k)$. Longer trips are selected for holding because this decision reduces the average remaining travel distance of vehicles circulating in the network, thereby improving overall vehicle exit rate. A detailed theoretical justification is provided in Section 2.2.
Note this paper assumes that the layout of parking resources, including the locations and parking capacities, $K^{p}$’s, is known. The first term in Eq. (6) is the holding demand, and the second term is the available parking supply for movement $(i,j,k)$, respectively.
@@
2.2 Theoretical justification for holding long-remaining-distance trips
The decision to hold vehicles with long remaining travel distances provides both direct and indirect performance benefits. The direct benefit arises from reducing network accumulation during congested conditions. Under the unimodal MFD relationship, decreasing accumulation in the congested regime increases network flow, thereby improving system performance. Beyond this accumulation effect, additional benefits emerge from selectively holding vehicles with large remaining travel distances. According to MFD theory, the average network flow can be expressed as the total vehicle-distance traveled per unit time:
$Q=\frac{\sum_{v\in\eta}s_{v}l_{v}}{L}$ (9)
where $\eta$ is the set of vehicles on the network, $s_{v}$ is the speed of vehicle $v$, $l_{v}$ is a characteristic link length, and $L$ is the total network length. Equivalently, drawing on Little’s Law, the average trip completion rate $\lambda$ relates to accumulation $n$ and mean remaining travel distance $\tilde{d}$ as:
$n=\lambda\frac{\tilde{d}}{\tilde{s}}$ (10)
or more compactly, since the average travel speed $\tilde{s}$ can be solely determined by accumulation, assuming traffic distribution is homogeneous, the trip completion rate can be expressed as:
$\lambda=\frac{Q(n)}{\tilde{d}}$ (11)
Equation 11 provides a key insight: for a given accumulation level $n$ (and thus a given flow $Q(n)$ in the congested regime), the trip completion rate $\lambda$ is inversely proportional to the mean remaining travel distance $\tilde{d}$. Selectively holding vehicles with long remaining travel distances reduces $\tilde{d}$ among vehicles circulating in the network. This increases the trip completion rate, which subsequently reduces accumulation and further increases $Q(n)$ through the MFD relationship. The proposed strategy creates a self-reinforcing recovery process: holding long-distance vehicles reduces $\tilde{d}$, which in turn increases trip completion rate, which further reduces n and increases $Q(n)$. Consequently, holding long-distance trips accelerates congestion dissipation more effectively than indiscriminate holding or holding short-distance trips. The performance of this design choice is demonstrated in Section 4.2.5.